# arXiv:2603.02863v1[math.LO] 03 Mar 2026

### v1

@@ Abstract.
Motivated by the results for Magic: The Gathering presented in [2] and [1], we study a computability problem related to optimal play in Yu-Gi-Oh! Trading Card Game, a popular card game developed and published by Konami. We show that the problem of establishing whether, from a given game state, a computable strategy is winning is undecidable. In particular, not only do we prove that the Halting Problem can be reduced to this problem, but also that the same holds for the Kleene’s $\mathcal{O}$, thereby demonstrating that this problem is actually $\Pi^{1}_{1}$-complete. We extend this last result to all strategies with a reduction on the set of countable well orders, a classic $\boldsymbol{\Pi}^{1}_{1}$-complete set. For these reductions we present two legal decks (according to the current Forbidden & Limited List of Yu-Gi-Oh! Trading Card Game) that can be used by the player who goes first to perform them.
@@ 1. Introduction
Despite the simplicity of the proof, these results are quite interesting from the point of view of game theory. In [1], the author remarks: “We have established that optimal play in Magic: The Gathering is at least as hard as $\emptyset^{(\omega)}$, the first time any real or even realistic game has been shown to be this complex.”. Our result surpasses the complexity result for Magic: The Gathering, as far as we know, optimal play in Yu-Gi-Oh! TCG is the most complex problem known in any real game. Nonetheless, we believe that our techniques could likely be adapted to show the same result for optimal play in Magic: The Gathering. It would be interesting to know whether the same holds for other card games (like Pokémon TGC).

### v2

# Deciding winning strategies in Yu-Gi-Oh! TCG is hard

@@ Abstract.
Motivated by the results for Magic: The Gathering presented in [2] and [1], we study a (different) computability problem about winning strategies in Yu-Gi-Oh! Trading Card Game, a popular card game developed and published by Konami. We show that the problem of establishing whether, from a given game state, a given computable strategy is winning is undecidable. In particular, not only do we prove that the Halting Problem can be reduced to this problem, but also that this problem is actually $\Pi^{1}_{1}$-complete. We extend this last result to all strategies with a reduction on the set of countable well orders, a classic $\boldsymbol{\Pi}^{1}_{1}$-complete set. For these reductions we present two legal decks (according to the current Forbidden & Limited List of Yu-Gi-Oh! Trading Card Game) that can be used by the player who goes first to perform them.
2020 Mathematics Subject Classification:
Primary 03D35 ; Secondary 91A44, 03E15
@@ 1. Introduction
We believe that our techniques could likely be adapted to show the same result about winning stetegies in Magic: The Gathering. It would be interesting to know whether the same holds for other card games (like Pokémon TGC).